5. Two identical conducting spheres hang from a light-weight string carry 10 nC of charges each maintain equilibrium state at a distance of 5 mm from each other while making 0-10o from the vertical line. Find the weight of the spheres.
Answer #1

drawing an fbd and doing force balance we get

vertical force balance

T*cos(\theta) = W {W is the weight}

horizontal force balance

F= T*sin(\theta)

and also

F= kq2/r2 = 9*109*(10*10-9)2/ (0.005)2

F = 0.036 N

dividing both equations

tan(\theta) = F/W

W = F/(tan(\theta))

putting the values

W=0.2041 N